MailSystemsandAddressing in4.2BSD EricAllman† Britton-Lee,Inc. 1919AddisonStreet,Suite105. Berkeley,California94704. eric@Berkeley.ARPA ucbvax!eric ABSTRACT Routingmailthroughaheterogeneousinternetpresentsmanynewproblems. Amongtheworstoftheseisthatofaddressmapping.Historically,thishas beenhandledonanadhocbasis.However,thisapproachhasbecomeunmanageableasinternetsgrow. Sendmailactsaunified“postoffice”towhichallmailcanbesubmitted.Addressinterpretationiscontrolledbyaproductionsystem,whichcanparseboth oldandnewformataddresses.Thenewformatis“domain-based,”aflexible techniquethatcanhandlemanycommonsituations.Sendmailisnotintended toperformuserinterfacefunctions. SendmailwillreplacedelivermailintheBerkeley4.2distribution.Severalmajorhostsarenoworwillsoonberunningsendmail.Thischangewillaffectany usersthatroutemailthroughasendmailgateway.Thechangesthatwillbeuservisibleareemphasized. Themailsystemtoappearin4.2BSDwillcontainanumberofchanges.Mostofthese changesarebasedonthereplacementofdelivermailwithanewmodulecalledsendmail.Sendmailimplementsageneralinternetworkmailroutingfacility,featuringaliasingandforwarding, automaticroutingtonetworkgateways,andflexibleconfiguration.Ofkeyinteresttothemail systemuserwillbethechangesinthenetworkaddressingstructure. Inasimplenetwork,eachnodehasanaddress,andresourcescanbeidentifiedwithahostresourcepair;inparticular,themailsystemcanrefertousersusingahost-usernamepair.Host namesandnumbershavetobeadministeredbyacentralauthority,butusernamescanbe assignedlocallytoeachhost. Inaninternet,multiplenetworkswithdifferentcharacteristicsandmanagementsmustcommunicate.Inparticular,thesyntaxandsemanticsofresourceidentificationchange.Certainspecialcasescanbehandledtriviallybyadhoctechniques,suchasprovidingnetworknamesthat appearlocaltohostsonothernetworks,aswiththeEthernetatXeroxPARC.However,thegeneralcaseisextremelycomplex.Forexample,somenetworksrequirethattheroutethemessage †AconsiderablepartofthisworkwasdonewhileundertheemployoftheINGRESProjectattheUniversityofCaliforniaatBerkeley. MailSystemsandAddressingin4.2BSD 1 MailSystemsandAddressingin4.2BSD 2 takesbeexplicitlyspecifiedbythesender,simplifyingthedatabaseupdateproblemsinceonly adjacenthostsmustbeenteredintothesystemtables,whileothersuselogicaladdressing,where thesenderspecifiesthelocationoftherecipientbutnothowtogetthere.Somenetworksusea left-associativesyntaxandothersusearight-associativesyntax,causingambiguityinmixed addresses. Internetstandardsseektoeliminatetheseproblems.Initially,theseproposedexpandingthe addresspairstoaddresstriples,consistingof{network,host,username}triples.Networknumbers mustbeuniversallyagreedupon,andhostscanbeassignedlocallyoneachnetwork.Theuserlevelpresentationwaschangedtoaddressdomains,comprisedofalocalresourceidentification andahierarchicaldomainspecificationwithacommonstaticroot.Thedomaintechniqueseparatestheissueofphysicalversuslogicaladdressing.Forexample,anaddressoftheform “eric@a.cc.berkeley.arpa”describesthelogicalorganizationoftheaddressspace(user“eric”on host“a”intheComputerCenteratBerkeley)butnotthephysicalnetworksused(forexample, thiscouldgooverdifferentnetworksdependingonwhether“a”wereonanethernetorastoreand-forwardnetwork). Sendmailisintendedtohelpbridgethegapbetweenthetotallyadhocworldofnetworks thatknownothingofeachotherandtheclean,tightly-coupledworldofuniquenetworknumbers. Itcanacceptoldarbitraryaddresssyntaxes,resolvingambiguitiesusingheuristicsspecifiedbythe systemadministrator,aswellasdomain-basedaddressing.Ithelpsguidetheconversionofmes-sageformatsbetweendisparatenetworks.Inshort,sendmailisdesignedtoassistagracefultran-sitiontoconsistentinternetworkaddressingschemes. Section1definessomeofthetermsfrequentlyleftfuzzywhenworkinginmailsystems.Section2discussesthedesigngoalsforsendmail.Insection3,thenewaddressformatsandbasic featuresofsendmailaredescribed.Section4discussessomeofthespecialproblemsoftheUUCP network.Thedifferencesbetweensendmailanddelivermailarepresentedinsection5. DISCLAIMER:Anumberofexamplesinthispaperusenamesofactualpeopleandorganizations.Thisisnotintendedtoimplyacommitmentorevenan intellectualagreementonthepartofthesepeopleororganizations.Inparticular,BellTelephoneLaboratories(BTL),DigitalEquipmentCorporation(DEC), LawrenceBerkeleyLaboratories(LBL),Britton-LeeIncorporated(BLI),and theUniversityofCaliforniaatBerkeleyarenotcommittedtoanyoftheseproposalsatthistime.Muchofthispaperrepresentsnomorethanthepersonal opinionsoftheauthor. 1.DEFINITIONS Therearefourbasicconceptsthatmustbeclearlydistinguishedwhendealingwithmail systems:theuser(ortheuser’sagent),theuser’sidentification,theuser’saddress,andthe route.Thesearedistinguishedprimarilybytheirpositionindependence. 1.1.UserandIdentification Theuseristhebeing(apersonorprogram)thatiscreatingorreceivingamessage. Anagentisanentityoperatingonbehalfoftheuser—suchasasecretarywhohandles mymail.oraprogramthatautomaticallyreturnsamessagesuchas“IamattheUNI-COMconference.” Theidentificationisthetagthatgoesalongwiththeparticularuser.Thistagis completelyindependentoflocation.Forexample,myidentificationisthestring“EricAllman,”andthisidentificationdoesnotchangewhetherIamlocatedatU.C.Berkeley,at Britton-Lee,oratascientificinstituteinAustria. Version1.3 USENIX−Jan83LastMod3/12/83 MailSystemsandAddressingin4.2BSD 3 Sincetheidentificationisfrequentlyambiguous(e.g.,therearetwo“RobertHenry”s atBerkeley)itiscommontoaddotherdisambiguatinginformationthatisnotstrictlypart oftheidentification(e.g.,Robert“CodeGenerator”HenryversusRobert“SystemAdministrator”Henry). 1.2.Address Theaddressspecifiesalocation.AsImovearound,myaddresschanges.Forexample,myaddressmightchangefrom“eric@Berkeley.ARPA”to“eric@bli.UUCP”or“allman@IIASA.Austria”dependingonmycurrentaffiliation. However,anaddressisindependentofthelocationofanyoneelse.Thatis,my addressremainsthesametoeveryonewhomightbesendingmemail.Forexample,apersonatMITandapersonatUSCcouldbothsendto“eric@Berkeley.ARPA”andhaveit arrivetothesamemailbox. Ideallya“whitepages”servicewouldbeprovidedtomapuseridentificationsinto addresses(forexample,see[Solomon81]).Currentlythisishandledbypassingaround scrapsofpaperorbycallingpeopleonthetelephonetofindouttheiraddress. 1.3.Route Whileanaddressspecifieswheretofindamailbox,aroutespecifieshowtofindthe mailbox.Specifically,itspecifiesapathfromsendertoreceiver.Assuch,therouteis potentiallydifferentforeverypairofpeopleintheelectronicuniverse. Normallytherouteishiddenfromtheuserbythesoftware.However,somenetworks puttheburdenofdeterminingtherouteontothesender.Althoughthissimplifiesthesoftware,italsogreatlyimpairstheusabilityformostusers.TheUUCPnetworkisanexampleofsuchanetwork. 2.DESIGNGOALS Designgoalsforsendmail1include: (1)Compatibilitywiththeexistingmailprograms,includingBellversion6mail,Bellversion7mail,BerkeleyMail[Shoens79],BerkNetmail[Schmidt79],andhopefullyUUCP mail[Nowitz78].ARPANETmail[Crocker82]wasalsorequired. (2)Reliability,inthesenseofguaranteeingthateverymessageiscorrectlydeliveredorat leastbroughttotheattentionofahumanforcorrectdisposal;nomessageshouldever becompletelylost.Thisgoalwasconsideredessentialbecauseoftheemphasisonmail inourenvironment.Ithasturnedouttobeoneofthehardestgoalstosatisfy,especiallyinthefaceofthemanyanomalousmessageformatsproducedbyvarious ARPANETsites.Forexample,certainsitesgenerateimproperlyformatedaddresses, occasionallycausingerror-messageloops.Somehostsuseblanksinnames,causing problemswithmailprogramsthatassumethatanaddressisoneword.Thesemantics ofsomefieldsareinterpretedslightlydifferentlybydifferentsites.Insummary,the obscurefeaturesoftheARPANETmailprotocolreallyareusedandaredifficultto support,butmustbesupported. (3)Existingsoftwaretodoactualdeliveryshouldbeusedwheneverpossible.Thisgoal derivesasmuchfrompoliticalandpracticalconsiderationsastechnical. (4)Easyexpansiontofairlycomplexenvironments,includingmultipleconnectionstoasinglenetworktype(suchaswithmultipleUUCPorEthernets).Thisgoalrequires 1Thissectionmakesnodistinctionbetweendelivermailandsendmail. Version1.3 USENIX−Jan83LastMod3/12/83 MailSystemsandAddressingin4.2BSD 4 considerationofthecontentsofanaddressaswellasitssyntaxinordertodetermine whichgatewaytouse. (5)Configurationinformationshouldnotbecompiledintothecode.Asinglecompiled programshouldbeabletorunasisatanysite(barringsuchbasicchangesastheCPU typeortheoperatingsystem).Wehavefoundthisseeminglyunimportantgoaltobe criticalinreallife.Besidesthesimpleproblemsthatoccurwhenanyprogramgets recompiledinadifferentenvironment,manysitesliketo“fiddle”withanythingthat theywillberecompilinganyway. (6)Sendmailmustbeabletoletvariousgroupsmaintaintheirownmailinglists,andlet individualsspecifytheirownforwarding,withoutmodifyingthesystemaliasfile. (7)Eachusershouldbeabletospecifywhichmailertoexecutetoprocessmailbeingdeliveredforhim.Thisfeatureallowsuserswhoareusingspecializedmailersthatuseadifferentformattobuildtheirenvironmentwithoutchangingthesystem,andfacilitates specializedfunctions(suchasreturningan“Iamonvacation”message). (8)Networktrafficshouldbeminimizedbybatchingaddressestoasinglehostwherepossible,withoutassistancefromtheuser. Thesegoalsmotivatedthearchitectureillustratedinfigure1.Theuserinteractswitha mailgeneratingandsendingprogram.Whenthemailiscreated,thegeneratorcallssendmail, whichroutesthemessagetothecorrectmailer(s).Sincesomeofthesendersmaybenetwork serversandsomeofthemailersmaybenetworkclients,sendmailmaybeusedasaninternet mailgateway. Figure1—SendmailSystemStructure. Version1.3 USENIX−Jan83LastMod3/12/83 MailSystemsandAddressingin4.2BSD 5 3.USAGE 3.1.AddressFormats Argumentsmaybeflagsoraddresses.Flagssetvariousprocessingoptions.Followingflagarguments,addressargumentsmaybegiven.Addressesfollowthesyntaxin RFC822[Crocker82]forARPANETaddressformats.Inbrief,theformatis: (1)Anythinginparenthesesisthrownaway(asacomment). (2)Anythinginanglebrackets(“<>”)ispreferredoveranythingelse.ThisruleimplementstheARPANETstandardthataddressesoftheform username<machine-address> willsendtotheelectronic“machine-address”ratherthanthehuman“username.” (3)Doublequotes(")quotephrases;backslashesquotecharacters.Backslashesare morepowerfulinthattheywillcauseotherwiseequivalentphrasestocomparedifferently—forexample,userand"user"areequivalent,but\userisdifferentfrom eitherofthem.Thismightbeusedtoavoidnormalaliasingorduplicatesuppressionalgorithms. Parentheses,anglebrackets,anddoublequotesmustbeproperlybalancedand nested.Therewritingrulescontrolremainingparsing2. Althougholdstyleaddressesarestillacceptedinmostcases,thepreferredaddress formatisbasedonARPANET-styledomain-basedaddresses[Su82a].Theseaddressesare basedonahierarchical,logicaldecompositionoftheaddressspace.TheaddressesarehierarchicalinasensesimilartotheU.S.postaladdresses:themessagesmayfirstberoutedto thecorrectstate,withnoinitialconsiderationofthecityorotheraddressingdetails.The addressesarelogicalinthateachstepinthehierarchycorrespondstoasetof“naming authorities”ratherthanaphysicalnetwork. Forexample,theaddress: eric@HostA.BigSite.ARPA wouldfirstlookupthedomainBigSiteinthenamespaceadministratedbyARPA.Aquery couldthenbesenttoBigSiteforinterpretationofHostA.Eventuallythemailwouldarrive atHostA,whichwouldthendofinaldeliverytouser“eric.” 3.2.MailtoFilesandPrograms Filesandprogramsarelegitimatemessagerecipients.Filesprovidearchivalstorage ofmessages,usefulforprojectadministrationandhistory.Programsareusefulasrecipientsinavarietyofsituations,forexample,tomaintainapublicrepositoryofsystemsmessages(suchastheBerkeleymsgsprogram). Anyaddresspassingthroughtheinitialparsingalgorithmasalocaladdress(i.e,not appearingtobeavalidaddressforanothermailer)isscannedfortwospecialcases.Ifprefixedbyaverticalbar(“|”)therestoftheaddressisprocessedasashellcommand.Ifthe usernamebeginswithaslashmark(“/”)thenameisusedasafilename,insteadofa loginname. 2Disclaimer:Somespecialprocessingisdoneafterrewritinglocalnames;seebelow. Version1.3 USENIX−Jan83LastMod3/12/83 MailSystemsandAddressingin4.2BSD 6 3.3.Aliasing,Forwarding,Inclusion Sendmailreroutesmailthreeways.Aliasingappliessystemwide.Forwardingallows eachusertorerouteincomingmaildestinedforthataccount.Inclusiondirectssendmailto readafileforalistofaddresses,andisnormallyusedinconjunctionwithaliasing. 3.3.1.Aliasing Aliasingmapslocaladdressestoaddresslistsusingasystem-widefile.Thisfileis hashedtospeedaccess.Onlyaddressesthatparseaslocalareallowedasaliases;this guaranteesauniquekey(sincetherearenonicknamesforthelocalhost). 3.3.2.Forwarding Afteraliasing,ifanrecipientaddressspecifiesalocalusersendmailsearchesfora “.forward”fileintherecipient’shomedirectory.Ifitexists,themessageisnotsentto thatuser,butrathertothelistofaddressesinthatfile.Oftenthislistwillcontainonly oneaddress,andthefeaturewillbeusedfornetworkmailforwarding. Forwardingalsopermitsausertospecifyaprivateincomingmailer.Forexample, forwardingto: "|/usr/local/newmailmyname" willuseadifferentincomingmailer. 3.3.3.Inclusion InclusionisspecifiedinRFC733[Crocker77]syntax: :Include:pathname Anaddressofthisformreadsthefilespecifiedbypathnameandsendstoalluserslisted inthatfile. Theintentisnottosupportdirectuseofthisfeature,butrathertousethisasa subsetofaliasing.Forexample,analiasoftheform: project::include:/usr/project/userlist isamethodoflettingaprojectmaintainamailinglistwithoutinteractionwiththesystemadministration,evenifthealiasfileisprotected. Itisnotnecessarytorebuildtheindexonthealiasdatabasewhena:include:list ischanged. 3.4.MessageCollection Onceallrecipientaddressesareparsedandverified,themessageiscollected.The messagecomesintwoparts:amessageheaderandamessagebody,separatedbyablank line.Thebodyisanuninterpretedsequenceoftextlines. Theheaderisformatedasaseriesoflinesoftheform field-name:field-value Field-valuecanbesplitacrosslinesbystartingthefollowinglineswithaspaceoratab. Someheaderfieldshavespecialinternalmeaning,andhaveappropriatespecialprocessing. Otherheadersaresimplypassedthrough.Someheaderfieldsmaybeaddedautomatically, suchastimestamps. 4.THEUUCPPROBLEM OfparticularinterestistheUUCPnetwork.TheexplicitroutingusedintheUUCP environmentcausesanumberofseriousproblems.First,givingoutanaddressisimpossible Version1.3 USENIX−Jan83LastMod3/12/83 MailSystemsandAddressingin4.2BSD 7 withoutknowingtheaddressofyourpotentialcorrespondent.Thisistypicallyhandledby specifyingtheaddressrelativetosome“well-known”host(e.g.,ucbvaxordecvax).Second,it isoftendifficulttocomputethesetofaddressestoreplytowithoutsomeknowledgeofthe topologyofthenetwork.Althoughitmaybeeasyforahumanbeingtodothisundermany circumstances,aprogramdoesnothaveequallysophisticatedheuristicsbuiltin.Third,certainaddresseswillbecomepainfullyandunnecessarilylong,aswhenamessageisrouted throughmanyhostsintheUSENET.Andfinally,certain“mixeddomain”addressesare impossibletoparseunambiguously—e.g., decvax!ucbvax!lbl-h!user@LBL-CSAM mighthavemanypossibleresolutions,dependingonwhetherthemessagewasfirstroutedto decvaxortoLBL-CSAM. Tosolvethisproblem,theUUCPsyntaxwouldhavetobechangedtouseaddresses ratherthanroutes.Forexample,theaddress“decvax!ucbvax!eric”mightbeexpressedas “eric@ucbvax.UUCP”(withthehopthroughdecvaximplied).Thisaddresswoulditselfbea domain-basedaddress;forexample,anaddressmightbeoftheform: mark@d.cbosg.btl.UUCP HostsoutsideofBellTelephoneLaboratorieswouldthenonlyneedtoknowhowtogettoa designatedBTLrelay,andtheBTLtopologywouldonlybemaintainedinsideBell. TherearethreemajorproblemsassociatedwithturningUUCPaddressesintosomething reasonable:definingthenamespace,creatingandpropagatingthenecessarysoftware,and buildingandmaintainingthedatabase. 4.1.DefiningtheNamespace PuttingallUUCPhostsintoaflatnamespace(e.g.,“...@host.UUCP”)isnotpracticalforanumberofreasons.First,withover1600sitesalready,and(withtheincreasing availabilityofinexpensivemicrocomputersandautodialers)severalthousandmorecoming withinafewyears,thedatabaseupdateproblemissimplyintractableifthenamespaceis flat.Second,therearealmostcertainlynameconflictstoday.Third,asthenumberofsites growthenamesbecomeeverlessmnemonic. Itseemsinevitablethattherebesomesortofnamingauthorityforthesetoftop levelnamesintheUUCPdomain,asunpleasantapossibilityasthatmayseem.Itwill simplynotbepossibletohaveonehostresolvingallnames.Itmayhoweverbepossibleto handlethisinafashionsimilartothatofassigningnamesofnewsgroupsinUSENET. However,itwillbeessentialtoencourageeveryonetobecomesubdomainsofanexisting domainwheneverpossible—eventhoughthiswillcertainlybruisesomeegos.Forexample,ifanewhostnamed“blid”weretobeaddedtotheUUCPnetwork,itwouldprobably actuallybeaddressedas“d.bli.UUCP”(i.e.,ashost“d”inthepseudo-domain“bli”rather thanashost“blid”intheUUCPdomain). 4.2.CreatingandPropagatingtheSoftware Thesoftwarerequiredtoimplementaconsistentnamespaceisrelativelytrivial.Two modulesareneeded,onetohandleincomingmailandonetohandleoutgoingmail. Theincomingmodulemustbepreparedtohandleeitheroldornewstyleaddresses. New-styleaddressescanbepassedthroughunchanged.Oldstyleaddressesmustbeturned intonewstyleaddresseswherepossible. Theoutgoingmoduleisslightlytrickier.Itmustdoadatabaselookupontherecipientaddresses(passedonthecommandline)todeterminewhathoststosendthemessage to.Ifthosehostsdonotacceptnew-styleaddresses,itmusttransformalladdressesinthe headerofthemessageintooldstyleusingthedatabaselookup. Version1.3 USENIX−Jan83LastMod3/12/83 MailSystemsandAddressingin4.2BSD 8 Bothofthesemodulesarestraightforwardexceptfortheissueofmodifyingthe header.Itseemsprudenttochooseoneformatforthemessageheaders.Foranumberof reasons,BerkeleyhaselectedtousetheARPANETprotocolsformessageformats.However,thisprotocolissomewhatdifficulttoparse. Propagationissomewhatmoredifficult.Therearealargenumberofhostsconnected toUUCPthatwillwanttoruncompletelystandardsystems(forverygoodreasons).The strategyisnottoconverttheentirenetwork—onlyenoughofititalleviatetheproblem. 4.3.BuildingandMaintainingtheDatabase Thisisbyfarthemostdifficultproblem.Aprototypeforthisdatabasealready exists,butitismaintainedbyhandanddoesnotpretendtobecomplete. Thisproblemwillbereducedconsiderablyifpeoplechoosetogrouptheirhostsinto subdomains.Thiswouldrequireaglobalupdateonlywhenanewtopleveldomainjoined thenetwork.Amessagetoahostinasubdomaincouldsimplyberoutedtoaknown domaingatewayforfurtherprocessing.Forexample,theaddress“eric@a.bli.UUCP”might beroutedtothe“bli”gatewayforredistribution;newhostscouldbeaddedwithinBLI withoutnotifyingtherestoftheworld.Ofcourse,otherhostscouldbenotifiedasanefficiencymeasure. Theremaybemorethanonedomaingateway.AdomainsuchasBTL,forinstance, mighthaveadozengatewaystotheoutsideworld;anon-BTLsitecouldchoosetheclosest gateway.Theonlyrestrictionwouldbethatallgatewaysmaintainaconsistentviewofthe domaintheyrepresent. 4.4.LogicalStructure Logically,domainsareorganizedintoatree.Thereneednotbeahostactuallyassociatedwitheachlevelinthetree—forexample,therewillbenohostassociatedwiththe name“UUCP.”Similarly,anorganizationmightgroupnamestogetherforadministrative reasons;forexample,thename CAD.research.BigCorp.UUCP mightnotactuallyhaveahostrepresenting“research.” However,itmayfrequentlybeconvenienttohaveahostorhoststhat“represent”a domain.Forexample,ifasinglehostexiststhatrepresentsBerkeley,thenmailfromoutsideBerkeleycanforwardmailtothathostforfurtherresolutionwithoutknowingBerkeley’s(rathervolatile)topology.Thisisnotunliketheoperationofthetelephonenetwork. Thismayalsobeusefulinsidecertainlargedomains.Forexample,atBerkeleyit maybepresumedthatmosthostsknowaboutotherhostsinsidetheBerkeleydomain.But iftheyprocessanaddressthatisunknown,theycanpassit“upstairs”forfurtherexamination.Thusasnewhostsareaddedonlyonehost(thedomainmaster)mustbeupdated immediately;otherhostscanbeupdatedasconvenient. Ideallythisnameresolutionprocesswouldbeperformedbyanameserver(e.g., [Su82b])toavoidunnecessarycopyingofthemessage.However,inabatchnetworksuchas UUCPthiscouldresultinunnecessarydelays. 5.COMPARISONWITHDELIVERMAIL Sendmailisanoutgrowthofdelivermail.Theprimarydifferencesare: (1)Configurationinformationisnotcompiledin.Thischangesimplifiesmanyoftheproblemsofmovingtoothermachines.Italsoallowseasydebuggingofnewmailers. Version1.3 USENIX−Jan83LastMod3/12/83 MailSystemsandAddressingin4.2BSD 9 (2)Addressparsingismoreflexible.Forexample,delivermailonlysupportedonegateway toanynetwork,whereassendmailcanbesensitivetohostnamesandreroutetodifferentgateways. (3)Forwardingand:include:featureseliminatetherequirementthatthesystemaliasfile bewritablebyanyuser(orthatanupdateprogrambewritten,orthatthesystem administrationmakeallchanges). (4)Sendmailsupportsmessagebatchingacrossnetworkswhenamessageisbeingsentto multiplerecipients. (5)Amailqueueisprovidedinsendmail.Mailthatcannotbedeliveredimmediatelybut canpotentiallybedeliveredlaterisstoredinthisqueueforalaterretry.Thequeue alsoprovidesabufferagainstsystemcrashes;afterthemessagehasbeencollectedit maybereliablyredeliveredevenifthesystemcrashesduringtheinitialdelivery. (6)Sendmailusesthenetworkingsupportprovidedby4.2BSDtoprovideadirectinterface networkssuchastheARPANETand/orEthernetusingSMTP(theSimpleMail TransferProtocol)overaTCP/IPconnection. Version1.3 USENIX−Jan83LastMod3/12/83 REFERENCES [Crocker77]Crocker,D.H.,Vittal,J.J.,Pogran,K.T.,andHenderson,D.A.Jr., StandardfortheFormatofARPANetworkTextMessages.RFC733, NIC41952.In[Feinler78].November1977. [Crocker82]Crocker,D.H.,StandardfortheFormatofArpaInternetTextMessages.RFC822.NetworkInformationCenter,SRIInternational, MenloPark,California.August1982. [Feinler78]Feinler,E.,andPostel,J.(eds.),ARPANETProtocolHandbook. NIC7104,NetworkInformationCenter,SRIInternational,Menlo Park,California.1978. [Nowitz78]Nowitz,D.A.,andLesk,M.E.,ADial-UpNetworkofUNIXSys-tems.BellLaboratories.InUNIXProgrammer’sManual,Seventh Edition,Volume2.August,1978. [Schmidt79]Schmidt,E.,AnIntroductiontotheBerkeleyNetwork.Universityof California,BerkeleyCalifornia.1979. [Shoens79]Shoens,K.,MailReferenceManual.UniversityofCalifornia,Berkeley.InUNIXProgrammer’sManual,SeventhEdition,Volume2C. December1979. [Solomon81]Solomon,M.,Landweber,L.,andNeuhengen,D.,TheDesignofthe CSNETNameServer.CS-DN-2.UniversityofWisconsin,Madison. October1981. [Su82a]Su,Zaw-Sing,andPostel,Jon,TheDomainNamingConventionfor InternetUserApplications.RFC819.NetworkInformationCenter, SRIInternational,MenloPark,California.August1982. [Su82b]Su,Zaw-Sing,ADistributedSystemforInternetNameService. RFC830.NetworkInformationCenter,SRIInternational,Menlo Park,California.October1982. MailSystemsandAddressingin4.2BSD 10